A certain volume of a gas at 298k is heated such that it's volume and pressure are now four times their original values . what Is the new temperature?


1
Expert's answer
2021-10-30T01:10:53-0400

P1V1/T1=4P1×4V1/T2

P1V1/298=4P1×4V1/T2

T2= 298×4P1×4V1/P1V1

= 298×16=4768K